OpenStudy (anonymous):

15. What is the equation of the ellipse with vertices's (5, 0), (-5, 0) and co-vertices's (0, 4), (0, -4)?

OpenStudy (anonymous):

The equation for an ellipse. \[\frac{ x^2 }{ a^2 }+\frac{y^2}{b^2}=1\] If the long axis is along the x axis, the half the length is a, and b is the short one. Questions?